how to get 16 by adding three consecutive numbers?
Answer by anjulasahay(30)

how to get 16 by adding three consecutive numbers?
ans :
let x be the 1st number of the three consecutive numbers.
then according to problem
x+(x+1)+(x+1+1) = 16
3x+3=16
or 3x=13
or x = 13/3
hence the three nos are
13/3,(13/3)+1),(13/3)+1+1)
13/3,16/3,19/3